what are the factors of 6r^2+3rs-18s^2
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

what are the factors of 6r^2+3rs-18s^2
divide by 3
2r^2+rs-6s^2
2r^2+4rs-3rs+6r^2=0
2r(r+4s)-3s(r+4s)=0
(2r-3s)(r+4s)

Answer by ikleyn(53937)

.
what are the factors of 6r^2+3rs-18s^2 ?
~~~~~~~~~~~~~~~~~~~~~~~~


        The solution in the post by @mananth is incorrect - simply ignore it.
        See below my correct solution.


    6r^2 + 3rs - 18s^2 = 

    = 3*(2r^2 + rs - 6s^2) =

    = 3*(2r^2 + 4rs - 3rs - 6s^2) =

    = 3*(2r(r+2s) - 3s(r+2s)) =

    = 3*(2r-3s)(r+2s).    ANSWER

Solved correctly.